# Comparative study of magnetic exchange parameters and magnon dispersions in NiO and MnO from first principles

### v1

@@ DISCUSSION
In conclusion, our comparative study provides a valuable contribution to the ongoing efforts aimed at assessing the accuracy of first-principles methods for describing magnetic interactions in transition-metal compounds. While our analysis focused on two prototypical systems, NiO and MnO, a broader benchmarking across a wider range of materials, with varying crystal structures and chemistries, is essential to draw more general conclusions. Each method considered in this work, $\Delta E$, MFT, and TDDFPT+$U$, has its own strengths and limitations in terms of accuracy, robustness, and computational cost. The $\Delta E$ and MFT approaches are computationally efficient, making them suitable for large or complex systems; however, their accuracy depends on the number of exchange parameters included (see Fig. S1 in the SI) and, in the case of MFT, on the implementation details. TDDFPT+$U$, while significantly more demanding computationally, offers an accurate and physically consistent framework that does not rely on an arbitrary selection of exchange parameters in the underlying model. While a single method can provide useful insights, predictive modeling of magnetic excitations benefits significantly from the use of multiple approaches. Whenever feasible, using multiple computational approaches provides a more reliable and comprehensive understanding, particularly when interpreting or guiding experimental investigations. Looking ahead, the continuous development and systematic benchmarking of these methods across diverse material classes will be crucial to advancing our ability to predict and understand magnetic excitations from first principles with high confidence.
METHODS
All calculations are performed using Quantum ESPRESSO  (v7.2) 118, 119, 120. Ground-state calculations are conducted with the PW code 118 using local spin-density approximation (LSDA) for the exchange-correlation functional including the Hubbard $U$ correction (LSDA+UU121, 122 . We use norm-conserving scalar-relativistic pseudopotentials 123 from the PseudoDojo library 124. The plane-wave expansion of the Kohn-Sham wavefunctions is carried out with an 80 Ry kinetic-energy cutoff, and a 320 Ry cutoff is used for the charge density and potentials. The BZ is sampled using a $\Gamma$-centered $12\times 12\times 12$ $\mathbf{k}$ points grid for the 4-atom rhombohedral unit cell, unless stated otherwise, and type II antiferromagnetic (AFII) ordering is used (see Fig. 1). Spin-orbit coupling is neglected. Geometry optimization is performed using the Broyden-Fletcher-Goldfarb-Shanno (BFGS) algorithm 125, with convergence criteria set to $10^{-8}$ Ry for total energy, $10^{-5}$ Ry/bohr for forces, and $0.01$ Kbar for pressure. All calculations are performed using the optimized LSDA+$U$ structural parameters. For NiO, the rhombohedral lattice parameter is $a=5.03$ Å and the rhombohedral angle is $\alpha=33.65^{\circ}$, while for MnO, $a=5.32$ Å and $\alpha=34.16^{\circ}$. The resulting magnetic moments and band gap values can be found in Ref. 39.
The Hubbard $U$ parameters are computed using DFPT 86, 88 [see Eq. (4) in the SI] as implemented in the HP code 126, with Löwdin-orthogonalized atomic orbitals for Hubbard projectors 127. We employ uniform $\Gamma$-centered $\mathbf{k}$ and $\mathbf{q}$ points grids of size $8\times 8\times 8$ and $4\times 4\times 4$, respectively, for the 4-atom AFII unit cell. The calculations use a kinetic-energy cutoff of 90 Ry for the Kohn-Sham wavefunctions and 360 Ry for the charge density and potentials, achieving an accuracy of approximately 0.01 eV for the Hubbard parameters. The $U$ parameters are computed iteratively in a self-consistent manner, as outlined in Ref. 87, and include Hubbard forces and stresses in LSDA+$U$ structural optimizations 128. The final computed values for $U$ are 6.26 eV for the Ni($3d$) and 4.29 eV for the Mn($3d$) states in NiO and MnO, respectively.
The magnon energies are computed using TDDFPT+$U$ 39, as implemented in a turboMAGNON code 129. This method is based on computing the dynamical spin susceptibility using the Liouville-Lanczos approach 71 [see Eqs. (21) and (23) in the SI]. We use adiabatic LSDA+$U$ (ALSDA+$U$). The 4-atom AFII unit cell is used and the BZ is sampled with a $\Gamma$-centered $12\times 12\times 12$ $\mathbf{k}$ points grid. The calculations employ the pseudo-Hermitian flavor of the Lanczos recursive algorithm 130, 15, which includes an extrapolation technique for the Lanczos coefficients 131. A total of 8000 Lanczos iterations are performed to achieve convergence in the TDDFPT+$U$ calculations. The TDDFT+$U$ exchange parameters are derived by fitting the TDDFPT+$U$ magnon dispersions using LSWT [see Eqs. (7)–(9) in the SI].
The exchange interaction parameters are computed using MFT [see Eqs. (18) in the SI] as implemented in the TB2J code (v0.7.7.2) 51. Calculations are performed with the 4-atom AFII unit cell. MLWFs are generated using the Wannier90 code (v3.1.0) 132, based on the LSDA+$U$ ground state. Specifically, we generate MLWFs for Mn($3d$), Mn($4s$), and O($2p$) states in MnO, as well as for Ni($3d$), Ni($4s$), and O($2p$) states in NiO, resulting in a total of 18 MLWFs out of 45 considered Kohn-Sham states for each material. We confirm that the Wannier centers for the Mn($3d$) and Ni($3d$) states are accurately centered on the atoms. For the ground-state calculations, the MLWFs generation, and the calculation of $J_{ij}$, the BZ is sampled using a $\Gamma$-centered $18\times 18\times 18$ $\mathbf{k}$ points grid. This sampling ensures that the calculation of the $J_{ij}$ parameters converges with an accuracy of 0.01 meV.
To compute the exchange interaction parameters using the $\Delta E$ method, we have developed a workflow within the AiiDA package 133, which will be described in detail elsewhere. We start with the AFII ground state in the LSDA+$U$ framework for the 4-atom unit cell. We then construct four supercells of size $4\times 4\times 4$, and sample the BZ using a $\Gamma$-centered $4\times 4\times 4$ $\mathbf{k}$ points grid, ensuring the convergence of the $J_{ij}$ parameters with an accuracy of 0.01 meV. In these supercells, we initialize the original AFII configuration and then consider a target pair of spins in the following configurations: $\uparrow\uparrow$ (state 1), $\uparrow\downarrow$ (state 2), $\downarrow\uparrow$ (state 3), and $\downarrow\downarrow$ (state 4). For each configuration, we compute the total energy and then evaluate the exchange interaction parameters [see Eq. (13) in the SI]. This procedure allows us to determine the exchange parameters $J_{ij}$ depending on the selected target pair of spins in the supercell.
The magnon dispersions based on LSWT, using the exchange interaction parameters reported in Table 1, are computed with the SpinW code (v3.1) 134. The fitting of all magnon dispersions in Fig. 2 is performed using the minimal model [see Eq. (2)] and the analytic LSWT expressions [see Eqs. (7)–(9) in the SI].
DATA AVAILABILITY
The data used to produce the results of this work will be available in the Materials Cloud Archive 135.
@@ ACKNOWLEDGEMENTS
F. dos S., L.B., N.M., and I.T. acknowledge support from the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (Grant No. 205602). I.T. acknowledges support from the Swiss National Science Foundation (Grant No. 200021-227641 and No. 200021-236507). L.B. acknowledges the Fellowship from the EPFL QSE Center “Many-body neural simulations of quantum materials” (Grant No. 10060). G.M. acknowledges support by the MUR through the PRIN project “q-LIMA” (Grant No. 2020JLZ52N). This work was supported by a grant from the Swiss National Supercomputing Centre (CSCS) under project “lp18” and “mr33”. G.M. acknowledges the computational resources have been provided by computing@unipi, a computing service provided by the University of Pisa, the HPC center of the Fondazione Istituto Italiano di Tecnologia (IIT, Genova). G.M. also acknowledges the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support, in particular through the “ISCRA C” projects “HP10CAVD6L”, “HP10CLNRW9”, and “HP10CY46PW”.
Author contributions statement
F. dos S.: Methodology, Software, Investigation, Writing - review & editing. L.B.: Methodology, Software, Investigation, Writing - review & editing. G.M.: Methodology, Software, Investigation, Writing - review & editing. R.M.: Methodology, Software, Investigation, Writing - review & editing. N.M.: Funding acquisition, Investigation, Writing - review & editing. I.T.: Conceptualization, Methodology, Software, Investigation, Project administration, Writing - original draft, Writing - review & editing.
Competing interests
The authors declare no competing interests.

### v2

@@ V Conclusions
Our comparative study provides a valuable contribution to the ongoing efforts aimed at assessing the accuracy of first-principles methods for describing magnetic interactions in transition-metal compounds. While our analysis focused on two prototypical systems, NiO and MnO, a broader benchmarking across a wider range of materials, with varying crystal structures and chemistries, is essential to draw more general conclusions. Each method considered in this work, $\Delta E$, MFT, and TDDFPT+$U$, has its own strengths and limitations in terms of accuracy, robustness, and computational cost. The $\Delta E$ and MFT approaches are less computationally expensive than TDDFPT+$U$, making them more suitable for large or complex systems; however, their accuracy depends on the number of exchange parameters included (see Fig. S1 in the SM [115]) and, in the case of MFT, on the implementation details. In contrast, TDDFPT+$U$, while significantly more expensive in terms of computational resources, offers an accurate and physically consistent framework that does not rely on an arbitrary selection of exchange parameters in the underlying model. Moreover, the computational complexity of the three methods, briefly summarized in Sec. S3 of the SM [115], differs significantly: The $\Delta E$ method relies on a series of standard self-consistent DFT+$U$ supercell calculations in which selected magnetic moments are flipped relative to the ground state; The MFT approach is a perturbative method based on the Green’s function formalism and, in certain implementations, additionally requires wannierization; TDDFPT+$U$ involves the explicit solution of linear-response equations, which substantially increases both the computational time and memory requirements. Moreover, extending TDDFPT to more advanced functionals (e.g., DFT+UU+$V$ or hybrid functionals) is theoretically and technically more involved. Therefore, while a single method can provide useful insights, predictive modeling of magnetic excitations benefits significantly from the use of multiple approaches. Whenever feasible, using multiple computational approaches provides a more reliable and comprehensive understanding, particularly when interpreting or guiding experimental investigations. Looking ahead, the continuous development and systematic benchmarking of these methods across diverse material classes will be crucial to advancing our ability to predict and understand magnetic excitations from first principles with high confidence.
@@ ACKNOWLEDGEMENTS
F. dos S. thanks Manuel dos Santos Dias for fruitful discussions. F. dos S., L.B., N.M., and I.T. acknowledge support from the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (Grant No. 205602). I.T. acknowledges support from the Swiss National Science Foundation (Grant No. 200021-227641 and No. 200021-236507). L.B. acknowledges the Fellowship from the EPFL QSE Center “Many-body neural simulations of quantum materials” (Grant No. 10060). G.M. acknowledges support by the MUR through the PRIN project “q-LIMA” (Grant No. 2020JLZ52N). This work was supported by a grant from the Swiss National Supercomputing Centre (CSCS) under project “lp18”, “s1326” and “mr33”. The computational resources have been provided by computing@unipi, a computing service provided by the University of Pisa, the HPC center of the Fondazione Istituto Italiano di Tecnologia (IIT, Genova). G.M. also acknowledges the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources and support, in particular through the “ISCRA C” projects “HP10CAVD6L”, “HP10CLNRW9”, and “HP10CY46PW”.